### v2

# Mingyu Liu1,2,∗  Jiuhe Shu1,∗  Hui Chen1  Zeju Li1  Canyu Zhao1 Jiange Yang3  Shenyuan Gao4  Hao Chen1,†  Chunhua Shen1,5,†

Abstract
A fundamental challenge in embodied intelligence is learning state representations that are both compact and expressive for world modeling and decision making, yet existing methods often remain either redundant or miss task-critical information. We propose an unsupervised approach that learns a highly compressed two-token state representation using a lightweight encoder and a pre-trained Diffusion Transformer (DiT) decoder, capitalizing on its strong generative prior. Our representation is efficient, interpretable, and integrates seamlessly into existing VLA-based models, improving performance by 11.6% on LIBERO and 31% in real-world task success rate with minimal inference overhead. More importantly, we find that the difference between these tokens, obtained via latent interpolation, naturally represents the motion, which can be further decoded into executable robot actions. This emergent capability reveals that our representation captures dynamics without explicit motion supervision. We name our method StaMo for its ability to learn generalizable robotic Motion from compact State representation, which is encoded from static images, challenging the prevalent dependence on learning robotic motions with complex temporal modeling and video data. Our learned representations also enhance policy co-training, outperforming prior methods by 10.4% with improved interpretability. Moreover, our approach scales effectively across diverse data sources, including real-world robot data, simulation, and human egocentric video. Project page at https://aim-uofa.github.io/StaMo/.
footnotetext: Equal contribution. Corresponding authors.
1 Introduction
“What we observe as static is merely dynamic equilibrium.”
— Richard Feynman, The Feynman Lectures on Physics
Figure 1: An overview of our StaMo framework. Our method efficiently compresses and encodes robotic visual representations, enabling the learning of a compact state representation. Motion naturally emerges as the difference between these states in the highly compressed token space. This approach facilitates efficient world modeling and demonstrates strong generalization, with the potential to scale up with more data.
Learning reusable and generalizable representations is a cornerstone of intelligent robotics systems. While visual representations in VLAs retain rich perceptual details for multimodal fusion, state representations for world modeling and intermediate reasoning serve a different purpose: they are designed to enable efficient future prediction and bridging visual planning with action execution. This role demands two key properties: First, it must be highly compact. Unlike rich visual embeddings used for perception, a state representation’s proximity to action generation demands a focus on actionable information to ensure efficient future prediction. Second, it must be expressive despite its compactness. Common low-dimensional approaches, such as trajectories [47] or flow fields [19, 50], often fail on this front; they capture basic motion but lack the semantic richness to encode goal states, interaction dynamics, or structured spatial relationships.
Building on these principles, we propose a compact state representation learned using a Diffusion Autoencoder fine-tuned on robotics data. Our method compresses an image into a highly compact latent state (as few as 2 tokens of 1024 dimensions). To ensure the representation is rich, we initialize its decoder from a powerful DiT model pretrained on internet data [18], reasoning that the ability to reconstruct pixels accurately requires an implicit understanding of key state information like robot poses and object interactions. This representation integrates seamlessly into existing VLM architectures, extensive experiments show that our method significantly enhances model performance with negligible overhead to inference speed, achieving improvements of + 11.6% in LIBERO and + 31% in real-world deployments.
Interestingly, from this compact state representation, we make an important discovery: robot motion (which is also known as latent action) naturally emerges from the state representation space. By simply performing a linear interpolation between the latent encodings of start and goal observations, we can generate smooth, plausible, and dynamically consistent motion trajectories. This observation provides an alternative to the dominant paradigm of learning latent actions from video data [53, 11, 52, 15]. While using video seems intuitive as action is inherently temporal, it presents significant challenges: it demands complex, computationally expensive temporal models and often produces ambiguous, coarse-grained actions. This is because high-variance motion within video clips causes models to learn averaged-out representations, making per-frame learning both inefficient and representationally flawed.
These challenges motivated us to revisit the fundamental motivation for using video data, if the objective is to capture latent actions through frame-wise changes, why should we commit to learning complex motion extractors from suboptimal state representations? We therefore raise a more fundamental question: rather than explicitly modeling motion from sequences, can we learn a sufficiently expressive state representation from individual frames such that the simple difference between two states naturally encapsulates a meaningful latent action? Our work demonstrates this is not only feasible but also highly advantageous. Compared to traditional video-based models, our state-centric approach is more training-efficient and avoids the representational ambiguity caused by motion variations in video. The emergent latent actions exhibit strong generalization and seamless sim-to-real transferability. We validate this effectiveness through extensive co-training experiments, where our latent actions achieve superior performance over traditional and more complex latent action learning approaches.
The captivating properties of state and motion ultimately form the foundation of our method, dubbed StaMo, which learns generalizable Motion from compact State representation. While the state is static, the motion is dynamic, a harmonious balance is elegantly achieved in our approach. We hope our method would shed new light on future research. Our contributions are summarized as follows:
  • We propose StaMo, a novel framework that encodes a compact state representation from static images, from which motion naturally emerges.
  • Our representation can be efficiently utilized for world modeling and serves as an intermediate representation that bridges vision-language models (VLMs) and action expert modules. It can be seamlessly integrated into existing VLM-based frameworks, delivering improved performance with minimal inference overhead.
  • Our motion extraction approach offers enhanced flexibility, strong generalization capability, and excellent transferability. It can be effectively utilized for co-training downstream models and facilitating goal-image planning tasks.
  • Comprehensive simulation and real-world experiments, along with extensive visualizations, demonstrate the effectiveness of our approach, which can be readily scaled up with additional data.
2 Related Works
2.1 Robot Representation Learning
Previous robotics research on visual representations has often faced a fundamental trade-off: methods excelling at motion representation, like latent actions [11, 53, 14, 6, 20], flow [19, 51], or trajectories [47], typically lack a rich, compact state representation. Conversely, approaches that encode detailed state information from raw images [46, 5] or dense features [55, 26, 34, 49, 32, 40] are often too high-dimensional and redundant to effectively represent motion via simple differences. StaMo, as shown in Figure 2, overcomes this dichotomy by achieving a unique balance. It learns to compress and encode a robot’s visual state into a highly compact token space that is both expressive enough for complex tasks and minimal enough that motion can be elegantly and powerfully represented as the difference between two states. This unified approach not only enables efficient world modeling but also demonstrates superior generalization and scalability, providing a robust foundation for future robotic systems.
2.2 World Modeling in VLA
Benefiting from powerful vision-language models (VLMs) [3, 16, 4] pre-trained on large-scale internet data, vision-language-action (VLA) models have demonstrated significant potential in generalizing across manipulation tasks. Typically, VLAs [8, 58, 13, 22, 25, 7, 48, 42, 45, 30, 43] map visual information and language instructions into the robot’s action space through end-to-end training. A natural extension is to endow the model with world modeling capabilities. Inspired by recent unified generative and comprehension models, a promising direction is to enable large models to reason with images—predicting future visual states while inferring actions or language, thereby creating a mutually reinforcing process. Several works [46, 55, 12, 27] have explored this idea. However, these approaches either require decoding full images during inference or rely on overly redundant state representations, limiting their generalization to novel scenes. In contrast, our method employs a more compact representation that improves model performance with minimal impact on inference speed. Furthermore, our representation exhibits strong generalization ability, enabling adaptation to unseen scenes without requiring further fine-tuning of the encoder.
2.3 Latent Action Learning
Robot learning scalability is constrained by scarce and heterogeneous robotic data, prompting the use of large-scale, action-free internet videos to instill foundational physical and operational knowledge for improved generalization and data efficiency [33]. Numerous studies have explored learning discrete [53, 15, 9, 11, 41, 10] or continuous [52, 20] latent actions from unlabeled data and have demonstrated their effectiveness across a wide range of downstream tasks. However, most of these approaches rely on carefully designed model architectures and depend on extracting frames from continuous videos, which introduces limitations such as sensitivity to frame sampling intervals, potential temporal biases, and poorly interpretable, blurry action representations. In contrast, our approach reveals that actionable representations can inherently emerge from large-scale natural image collections—challenging the conventional belief that latent actions must be learned from video sequences. We further demonstrate that such image-emergent actions exhibit superior interpretability and stronger generalization capabilities.
Figure 2: Where is StaMo? This figure visualizes how different robotic representations fall on the spectrum of expressiveness versus compactness. StaMo uniquely occupies the ideal position, offering both a rich, expressive state representation and the ability to model motion from a highly compact space.
3 Method
3.1 Compress Image into Compact State
[…]
To this end, as illustrated in Figure 1, the encoder of our Diffusion Autoencoder, denoted as $\mathcal{E}$, consists of a pretrained DINOv2 model [36] serving as a frozen feature extractor, which is followed by a transformer-based compressor, to map the input observations into a sequence of compact token representations. The decoder $\mathcal{D}$, is a Diffusion Transformer (DiT) [37] conditioned on these tokens to reconstruct the original image observations. Our implementation is built upon the Stable Diffusion 3, where only the transformer-based compressor and the DiT decoder are trained. The DINOv2 encoder remains frozen throughout the training process. We employ the same Flow Matching [2, 31, 28] objective used in base model training to optimize the model, which minimizes the loss between the velocity predicted by $\mathcal{D}$ and the target velocity $u$:
$\mathbf{z}_{0}$ $=\tau(\mathbf{x}_{0}),$ (1)
${L_{DAE}}={\mathbb{E}}_{\mathbf{z}_{0},t}\|\mathcal{D}$ $(\mathbf{z}_{t},\mathcal{E}(\mathbf{x}_{0}),t)-u(\mathbf{z}_{t})\|^{2}_{2},$
where $\mathbf{x}_{0}$ is the input image observation and $\tau$ is the VAE encoder of pretrained diffusion model to produce latent $\mathbf{z}_{0}$. $\mathbf{z}_{t}=(1-\sigma_{t})\mathbf{z}_{0}+\sigma_{t}\epsilon$ is the linear interpolation between pure noise $\epsilon$ and $\mathbf{z}_{0}$. After training, The trained encoder $\mathcal{E}$ is able to yield the state representation $s$.
3.2 Interpolating Motion from States
A key advantage of our learned representation is its capacity to serve a dual role, representing both state and motion. This duality resolves a fundamental trade-off inherent in conventional robotics representations, which we analyze comparatively in Figure 2. On one hand, traditional motion representations such as end-effector poses, optical flow, trajectories, or recent latent action models, are typically low-dimensional and compact. Their strength lies in capturing dynamics, often through simple differencing, making them effective for representing actions. However, this compactness comes at the cost of expressivity: they lack the rich visual context necessary to reconstruct a plausible or detailed state. On the other hand, prevailing state representations, derived from encoding raw image observations or supplementary inputs like depth and segmentation maps, capture rich semantic and geometric details. While highly expressive, these high-dimensional features are computationally burdensome and fail to intrinsically encode the dynamics or motion required to transition between states. Our approach elegantly bridges this gap. By defining motion as the vector difference between consecutive compact state tokens $a_{t}=s_{t+1}-s_{t}$, we achieve a representation that is both compact and expressive, unifying the description of “what the world looks like” and “how the world changes”. The effectiveness of representing motions as the differences between states is demonstrated in our experiments in Section 4.6.
3.3 StaMo for Efficient World Modeling
Existing VLA models, such as our OpenVLA baseline, typically operate as reactive policies. They learn a direct mapping from the current visuolinguistic context to a sequence of low-level actions (e.g., 7-DoF end-effector controls). While effective, this paradigm does not explicitly compel the agent to reason about the physical consequences of its actions or how the world will change as a result. We hypothesize that endowing the agent with a predictive world model to anticipate future visual states will improve the future-planning capability of model. This auxiliary task of predicting “what happens next” should, in turn, regularize the policy and improve the quality of the primary action prediction task.
The compact state representations produced by our StaMo encoder $\mathcal{E}$ are ideally suited for this purpose, which provide a concise yet rich summary of the environment for the model to reason about the future. To this end, we integrate the compact representations of StaMo into the OpenVLA architecture and train the model to jointly predict the next state and the corresponding action. Technically, we achieve this by attaching a lightweight MLP head to OpenVLA’s autoregressive backbone, which is tasked specifically with predicting the subsequent state representation. The model is optimized with a composite loss function that balances action generation and future-state prediction:
$L_{total}=\lambda_{action}L_{action}+\lambda_{future}(L_{mse}(s_{pred},s_{gt})\\ +L_{1}(s_{pred},s_{gt})),$ (2)
where $L_{action}$ is the standard cross-entropy loss for next-token prediction, anchoring the model to its primary control task. The world model objective combines Mean Squared Error (MSE) and L1 losses to enforce accurate regression of the ground-truth future state. In our experiments, we set $\lambda_{action}=\lambda_{future}$, reflecting our hypothesis that learning to predict is as crucial as learning to act. Our experimental results validate this hypothesis, demonstrating that compelling the VLA to predict future states significantly improves task success rates, as detailed in Section 4.2.
3.4 StaMo for Latent Motions Co-Training
While Section 4.1 provides a qualitative visualization of the emergent motion, the abstract nature of our latent motion representation makes direct quantitative comparison challenging. To rigorously assess its effectiveness, we employ a policy co-training strategy. This approach evaluates the utility of our motion representation on downstream tasks by jointly training a policy model on a small set of action-labeled robot data alongside a larger set of action-less video data. We infer a latent motion, $m_{t}$, for each pair of consecutive video frames by calculating the difference between their compact state representations encoded by our model $E(\cdot):m_{t}=E(o_{t+1})-E(o_{t})$, This process generates pseudo-action labels for the video data, creating a unified dataset where our emergent motions and ground-truth robot actions are learned jointly within a single policy model. This approach unlocks the potential to learn from vast and diverse video sources without requiring explicit action labels.
We tested this co-training framework in Section 4.5, our method leads to substantial performance gains compared with previous method. The results confirm that the motions emerging from StaMo serve as an effective and robust proxy for true actions, providing a scalable pathway to enhance policy learning by leveraging the wealth of unlabeled visual data.
4 Experiments
4.1 Qualitative Analysis of StaMo
StaMo demonstrates the ability to perform high-quality reconstructions of robotic manipulation images with strong generalization capabilities. It achieves excellent results on both in-domain and out-of-domain data. Quantitative reconstruction results are shown in Table 1, and some qualitative results are presented in Figure 3. More results can be found in Appendix. For main experiments, we use DROID [23] and LIBERO [29] as our training data.
Table 1: Reconstruct Performance comparison of different datasets using our StaMo encoder.
Dataset libero_10 libero_90 libero_goal libero_object libero_spatial Droid Maniskill(OOD)
PSNR (dB) 25.5194 27.2470 24.6467 27.0011 25.9683 20.2492 22.1673
SSIM 0.8909 0.8962 0.8926 0.9105 0.8984 0.7346 0.8824
Figure 3 also showcases StaMo’s powerful motion interpolation ability. By linearly interpolating the tokens in the latent space, the decoded images exhibit excellent continuity and plausible motion. Furthermore, we demonstrate the results of motion transfer. By taking the difference between tokens in the latent space, the resulting latent motion proves to be highly transferable—both in sim-to-sim, sim-to-real, and real-to-sim scenarios. This indicates that the learned motion is not scene-specific but possesses strong generalization capabilities.
Figure 3: Reconstruct images using our StaMo encoder with as few as two 1024-dimensional tokens. The first row shows the ground truth, and the second row shows the predicted results, with corresponding PSNR and SSIM metrics listed below. The results demonstrate that StaMo can preserve high image fidelity and structural similarity even under extremely compressed state representations.
4.2 World Modeling Results
In this section, we evaluate the application of our StaMo representation for efficient world modeling. A key advantage of our approach is its ability to integrate seamlessly into existing VLM-based frameworks. We demonstrate this by conducting experiments on two strong baselines: OpenVLA and its successor, OpenVLA-OFT. We specifically chose OpenVLA-OFT because it introduces architectural improvements like parallel decoding and action chunking, which result in a significantly longer action horizon. This requires the model to predict a state representation at a more distant future step, making it a more challenging and relevant benchmark for world modeling.
To validate our approach, we retrained both OpenVLA and OpenVLA-OFT, replacing their original vision encoders with our StaMo encoder and training them to predict StaMo states. We denote these modified architectures as OpenVLA* and OpenVLA-OFT*; the original, unmodified models are presented without the asterisk. For our OpenVLA-OFT* variant, we introduced a minor architectural modification, using only the third-person camera view and removing the wrist camera input. We evaluate the performance of our StaMo representation on the LIBERO benchmark to demonstrate its superior capability in assisting VLA manipulation through enhanced world modeling. For each task, we conduct 1,000 evaluation rollouts, and all results are presented as percentage success rates.
Table 2: Performance comparison of different methods (%).
Method Spatial Object Goal Long Average
DP [17] 78.3% 92.5% 68.3% 50.5% 72.4%
Octo-Base [42] 78.9% 85.7% 84.6% 51.1% 75.1%
Spatial-VLA [39] 88.2% 89.9% 78.6% 55.5% 78.1%
CoT-VLA [56] 87.5% 91.6% 87.6% 69.0% 83.9%
$\pi$0-FAST [38] 96.4% 96.8% 88.6% 60.2% 85.5%
WorldVLA [12] 87.6% 96.2% 83.4% 60.0% 81.8%
UniVLA [46] 95.4% 98.8% 93.6% 94.0% 96.5%
OpenVLA [25] 84.7% 88.4% 79.2% 53.7% 76.5%
OpenVLA* + DINOv2 Feature 88.6% 90.4% 83.5% 61.4% 80.9%
OpenVLA* + StaMo state 92.3% 94.8% 88.1% 75.2% 87.6%
OpenVLA* + StaMo motion 93.1% 95.1% 87.4% 76.9% 88.1%
OpenVLA-OFT [24] 93.7% 94.2% 89.7% 91.3% 92.2%
OpenVLA-OFT* + DINOv2 Feature 94.1% 95.0% 91.2% 93.0% 93.3%
OpenVLA-OFT* + StaMo state 96.8% 98.9% 95.0% 96.3% 96.8%
OpenVLA-OFT* + StaMo motion 95.3% 97.1% 94.1% 92.9% 94.9%
Both the StaMo state and motion representations can serve as predictive targets for a world model, and their performance is detailed in Table 2. Our findings indicate that the optimal choice of representation depends on the policy’s action horizon. For the standard OpenVLA, which predicts actions over a short, single-step interval, the motion representation yields better performance. This is because the latent motion is conceptually analogous to the delta end-effector (EE) pose used in frameworks like LIBERO, providing a direct and incremental target. Conversely, for OpenVLA-OFT, which plans over a longer execution step, the state representation is recommended, as it acts as a stable goal-conditioning signal.
4.3 Inference Speed Comparison
A key advantage of StaMo lies in its utilization of two highly compressed tokens to assist VLM-based models with world modeling. During the inference stage, it obviates the need to decode full images; the model only needs to predict the tokens themselves. This approach significantly enhances the efficiency of the world modeling process. We conducted a comparative analysis of the model’s inference frequency before and after integrating the StaMo representation, and also benchmarked its speed against other world modeling methods. Our findings indicate that StaMo introduces significantly less inference overhead. The comparison of different model’s inference frequency is presented in Table 3.
Table 3: Inference Speed Comparison.The table presents a comparison of inference speeds for different methods, with all models deployed on a single RTX 4090 GPU. We benchmark against other world model-based VLA approaches, such as UniVLA [46] and WorldVLA [12]. As the results show, our method achieves a higher inference frequency compared to previous methods, including directly using DINOv2 features.Our approach introduces minimal computational overhead compared to the original baseline model.
Method Frequency (Hz) $\uparrow$
WorldVLA [12] 2.27
UniVLA [46] 2.65
OpenVLA [25] 4.16
OpenVLA + DINOv2 Feature 2.86
OpenVLA + StaMo 4.02
OpenVLA-OFT [24] 18.24
OpenVLA-OFT + DINOv2 Feature 11.42
OpenVLA-OFT + StaMo 17.82
Table 4: Zero-Shot Performance on Maniskill [21].
Method Pick Cube Push Cube Stack Cube Pull Cube Tool
OpenVLA [25] 28% 4% 18% 16%
UniVLA [46] 32% 7% 22% 20%
WorldVLA [12] 30% 10% 17% 22%
OpenVLA + StaMo state 41% 32% 48% 36%
4.4 Zero-Shot Transfer to New Environment
Due to the compact nature of the StaMo representation, which focuses on manipulation-relevant information rather than extraneous visual details, our model exhibits superior generalization capabilities. We validated this in Table 4 by taking the model trained on the LIBERO benchmark and applying it directly to the ManiSkill [21] environment. Crucially, all methods were deployed in this zero-shot transfer setting. We evaluated each task over 100 trials, with results presented as percentage success rates. The outcomes demonstrate that our approach achieves stronger performance than competing models under these conditions.
4.5 Policy Co-training Results
In Table 5, we detail the results of our policy co-training. Our approach utilizes a DIT-based model on the RDT architecture, trained on a per-task dataset of 10 robot trajectories and 40 video trajectories with StaMo-generated pseudo-actions. We evaluated the final model checkpoint for each task across 20 trials, averaging the success rate over three full runs. StaMo significantly outperforms both LAPA [53] and ATM [47], suggesting that its pseudo-action labels are more faithful to the original actions.
Table 5: Performance comparison of RDT with different data configurations (%).
Method Spatial Object Goal Long Average
RDT (all Real) 91.6% 93.3% 86.7% 73.3% 86.2%
RDT (1 Real) 71.7% 70.0% 66.7% 43.3% 62.9%
RDT (1Real+ 4ATM [47]) 83.3% 81.7% 71.7% 56.7% 73.4%
RDT (1Real+ 4LAPA [53]) 80.0% 76.7% 75.0% 65.0% 74.2%
RDT (1Real + 4StaMo) 90.0% 91.6% 86.7% 70.0% 84.6%
4.6 Action Linear Probing Results
A central claim of our work is that an effective latent action can be formulated as the simple vector difference between two latent states encoded by StaMo. To quantitatively validate the quality and utility of these emergent latent actions, we employ a linear probing protocol [1], a standard methodology for evaluating the efficacy of learned representations, to measure how much explicit, predictable information about the motion is contained within our latent action formulation.
We first construct a dataset by randomly sampling tuples from a large collection of robot trajectories. Each sampled tuple is of the form $\left(\mathbf{I}_{n},\mathbf{I}_{n+k},\mathbf{A}_{n}\right)$, where $\mathbf{I}_{n}$ is a starting image, $\mathbf{I}_{n+k}$ is the goal image after a horizon of $k$ steps, and $\mathbf{A}_{n}=\left(\mathbf{a}_{n},…,\mathbf{a}_{n+k-1}\right)$ is the sequence of ground-truth robot actions executed between them. Each image pair is then encoder into latent action, $\Delta z$, using our frozen StaMo encoder, $\mathcal{E}$:
$\Delta z=\mathcal{E}\left(\mathbf{I}_{n+k}\right)-\mathcal{E}\left(\mathbf{I}_{n}\right).$ (3)
We then train a lightweight Multi-Layer Perceptron (MLP) to predict the action sequence $\mathbf{A}_{n}$ from the latent action representation $\Delta z$. The performance is measured by the Mean Squared Error (MSE) between the predicted action sequence $\hat{\mathbf{A}}_{n}$ and the ground-truth sequence $\mathbf{A}_{n}$. This process can be formally described as follows.
$\operatorname{MSE}\left(\mathbf{A}_{n},\mathbf{\hat{A}}_{n}\right)=\operatorname{MSE}\left(\mathbf{A}_{n},\operatorname{MLP}(\Delta z)\right).$ (4)
A low LP-MSE score provides strong quantitative evidence that the vector difference between our latent states serves as a highly informative and linearly separable representation for robotic motions.
Figure 4: Linear Probing MSE results. We compare our method against three baselines. Our method consistently achieves the lowest MSE across all horizons.
To validate our approach, we benchmark our latent action representation against several baselines. The first set of baselines tests the importance of our structured latent space against alternatives operating on raw data: Pooled Delta Image (using pixel-wise differences) and Delta DINOv2 Features (using feature-space differences). The second baseline, LAPA [54], contrasts our deterministic state-difference formulation with a state-of-the-art autoregressive model that generates latent action tokens.
For a fair comparison, all representations are used to train an identical lightweight MLP. We evaluate non-generative methods across 1, 2, 4, and 8-step prediction horizons, while LAPA is assessed at its native 1-step horizon. Once trained, the MLP can directly map a current observation and a goal image to an executable action trajectory on the test set. The results in Figure 4 confirm the superiority of our method.
4.7 Goal-Conditioned Task Planning
Once the linear head is trained, it becomes capable of translating abstract representations into concrete, executable actions. We evaluated the goal-conditioned task planning capabilities of StaMo and DINO-WM [57] using the ManiSkill benchmark. ManiSkill was specifically selected because it was not included in the pre-training datasets for either StaMo or DINO-WM. For this comparison, we trained only the linear head while keeping the encoders of both StaMo and DINO-WM frozen. Table 6 presents the performance comparison, and further implementation details are available in the appendix. We evaluated each task over 100 trials, with results presented as percentage success rates.
Table 6: Goal-Conditioned Task Planning.
Method Pick Cube Push Cube Stack Cube Pull Cube Tool
DINO-WM [57] 12% 8% 4% 18%
StaMo 22% 32% 16% 38%
4.8 Real World Experiments
Table 7: Performance comparison on short and long horizon tasks.Best results are in bold and second-best results are underlined.
Short horizon tasks Long horizon tasks
Method
Pick up
the [toyname]
put the toy
into the basket
Open the
drawer
Average
put all cups
into the basket
put the toy
into the drawer
Stack all
cups
Average
All
Average
Spatial-VLA [39] 0.45 0.50 0.20 0.38 0.35 0.20 0.20 0.25 0.32
WorldVLA [12] 0.55 0.20 0.15 0.30 0.25 0.40 0.15 0.26 0.28
UniVLA [46] 0.55 0.30 0.35 0.40 0.35 0.30 0.10 0.25 0.33
OpenVLA [25] 0.35 0.30 0.25 0.30 0.20 0.30 0.15 0.20 0.25
OpenVLA + StaMo state 0.60 0.55 0.50 0.60 0.55 0.55 0.45 0.52 0.56
OpenVLA-OFT + StaMo state 0.70 0.65 0.60 0.65 0.70 0.65 0.55 0.63 0.64
Experiment Setup Our real-world experimental benchmark comprises six manipulation tasks (three short-horizon and three long-horizon) to comprehensively evaluate the effectiveness of our StaMo representation in facilitating world modeling and decision-making across a spectrum of task complexities, as shown in Figure 5. For each task, we collected 50 human demonstrations and evaluated performance over 20 trials. The results, presented as percentage success rates, are shown in Table 7.
Figure 5: Real World Setting and Tasks. We designed a benchmark of six real-world robotics tasks, spanning both short- and long-horizon challenges, to evaluate the effectiveness of our representation for learning world models.
4.9 Ablation Study
4.9.1 Ablation of Token Dimension
We ablate the hidden dimension of StaMo’s 2-token latent state (trained on LIBERO datasets, evaluated via PSNR/SSIM). As shown in Table 8, across 256/512/1024 dimensions. This indicates hidden dimension has limited impact on performance, as the pre-trained DiT decoder’s generative prior offsets potential information loss in lower dimensions.
Table 8: Reconstruction performance with 2 tokens under different hidden dimensions using our StaMo encoder.
hidden_dim Datasets libero_10 libero_90 libero_goal libero_object libero_spatial
256 PSNR (dB) 26.5969 28.5858 27.5070 31.2779 25.7993
SSIM 0.9082 0.8948 0.9327 0.9478 0.9118
512 PSNR (dB) 27.3322 29.5773 26.2000 31.0129 24.6279
SSIM 0.9180 0.9127 0.9194 0.9444 0.9009
1024 PSNR (dB) 27.0418 29.4810 26.4807 31.2653 25.9964
SSIM 0.9186 0.9124 0.9194 0.9480 0.9155
4.9.2 Ablation of Additional Training Data
A natural question following our initial validation is the scalability of our approach. To explore this, we investigated the effect of the StaMo training data volume and diversity on real-world task performance. We compared several data configurations: (1) simulation data only, (2) simulation + DROID [23] data (our main setting), (3) simulation + DROID + OXE [35] data, and (4) simulation + DROID + OXE + egocentric human (Ego) data [44]. The results, clearly depicted in Figure 6, confirm that our model is highly scalable. This verifies that incorporating larger volumes of data and greater task diversity, such as varied robot embodiments from OXE and human-centric data, steadily improves StaMo’s performance capabilities.
Figure 6: Scaling Performance. Performance of our model can be scale with more data, including human ego-centric data.“S” means Simulation data only,“D” represents DROID [23] data, “O” represents OXE [35] data, “Ego” means Ego-centric data.
4.9.3 Ablation of Pretrained Diffusion Decoder
We compared the performance of using a pre-trained diffusion decoder against training one entirely from scratch. The decoder trained from scratch required a significantly longer time to converge and ultimately was unable to match the same level of performance as its pre-trained counterpart. As shown in Table 9. This finding demonstrates the critical importance of pre-training on large-scale natural image datasets. Such pre-training provides a crucial prior, which is highly beneficial for learning a more effective representation of the robotics data distribution.
Table 9: Performance comparison between using pretrained diffusion decoder and training from scratch.
Method Spatial Object Goal Long Average
Train from Scratch 90.3% 92.7% 86.5% 73.4% 85.7%
w. Pretrain Diffusion Decoder 92.3% 94.8% 88.1% 75.2% 87.6%
5 Conclusion
In this work, we introduced StaMo, a framework that extracts reusable latent action vectors from static images via a transformer-based tokenizer and a diffusion decoder. We demonstrate that linear interpolation in this latent space produces smooth, physically plausible motions, and that these vectors generalize zero-shot to unseen senarios, suggesting that large-scale visual models implicitly learn a linearized dynamics manifold. Our approach offers a pathway toward scalable unsupervised skill discovery from diverse visual data, bridging the gap between static perception and dynamic action.
Acknowledgments
This work was supported in part by the Zhejiang Provincial Pioneering Science and Technology Program (2025), and by the National Natural Science Foundation of China (No. 62576315).